# Image Generation from Contextually-Contradictory Prompts

### v1

@@ Abstract.
Figure 1. Our method addresses contextual contradictions in text-to-image generation. These contradictions arise when one concept implicitly conflicts with another due to the model’s learned associations. For example, if a concept like “nest” is strongly entangled with “birds”, it may conflict with another concept in the prompt such as “rooster”, leading the model to ignore or distort part of the semantic meaning.
1. Introduction
@@
Text-to-image diffusion models have demonstrated remarkable capabilities in generating high-quality and diverse visual content from natural language prompts (Rombach et al., 2022; Ramesh et al., 2021; Ho et al., 2020). However, achieving precise semantic alignment between the generated image and the conditioning prompt remains an open challenge. This issue becomes particularly pronounced when the target prompt lies outside the model’s training distribution, such as prompts that combine semantically plausible yet statistically uncommon or unconventional concepts. For example, as shown in Figure 1, generating an image from the prompt “a rooster in a nest” often results in colorful birds sitting in a nest. This is due to the model’s strong prior entanglement between bird species and nests, which implicitly contradicts the notion of a rooster.
@@
We refer to this phenomenon as Contextual Contradiction, a conflict between two concepts that arises not from direct semantic opposition, but from the model’s associations learned during training. More precisely, we say that concept $A$ contextually contradicts concept $B$ if the model’s prior entangles $A$ with concept $C$, and $B$ contradicts $C$. In Figure 1, we illustrate this with a blowing dragon, which stands in contextual contradiction with the water due to its entanglement with fire.
The phenomenon of contextual contradiction in text-to-image models relates to the broader issue of spurious correlations in deep learning. Models often rely on shortcuts, which are predictive patterns in the training data that do not reflect true semantics (Geirhos et al., 2020). In NLP, models exploit syntactic heuristics rather than reasoning (McCoy et al., 2019), while in computer vision, they may depend on background cues rather than object features (Xiao et al., 2021; Singh et al., 2020). For example, classifiers may fail to recognize cows on beaches due to background biases (Beery et al., 2018). In this paper, we identify similar biases in generative models, where contextual contradictions arise when prompts include plausible but uncommon concept combinations.
@@
To address this issue, we build on the observation that the denoising process follows a coarse-to-fine progression, during which different semantic attributes (e.g., background, pose, shape, and texture) emerge at distinct stages (Chefer et al., 2023; Balaji et al., 2022; Patashnik et al., 2023). Our key idea is to guide the model at each stage of denoising with the information most relevant to the type of content being formed at that point. To achieve this, we decompose the original prompt into a sequence of proxy prompts, each aligned with the attributes expected at a specific stage and designed to avoid contextual contradictions.
@@
Ensuring that proxy prompts preserve the original intent while avoiding contextual contradictions requires a broad understanding of the real world. For example, it involves understanding that a bear is entangled with specific poses, such as walking on all fours or standing upright, which in turn, contradicts the handstand pose. Our method leverages a large language model (LLM) to analyze the target prompt, identify contextual contradictions, and construct suitable proxy prompts.
@@
Through extensive experiments, we demonstrate the effectiveness of our method in generating images from contextually contradicting text prompts. By introducing prompt information at targeted stages, our approach generates precise combination of semantic attributes while avoiding undesired entanglement of unrelated attributes. Compared to previous methods, our stage-dependent prompt decomposition leads to more faithful and semantically aligned generations.
2. Related Work
Learned Spurious Correlations
Machine learning models are known to be sensitive to spurious correlations in their training data (Geirhos et al., 2020; McCoy et al., 2019; Ye et al., 2024). This sensitivity often leads to a significant drop in performance when training-time associations between concepts do not hold at test time. Prior work has extensively studied this issue in discriminative computer vision tasks, showing, for example, that recognition models tend to rely heavily on background cues (Singh et al., 2020; Xiao et al., 2021; Beery et al., 2018).
@@ 2. Related Work
In our work, we show that text-to-image models exhibit similar behavior. When given contextually contradictory prompts – combinations of concepts that conflict with correlations seen during training – diffusion models often fail to generate images that accurately reflect the prompt. We evaluate this behavior using the Whoops! dataset (Bitton-Guetta et al., 2023), which contains prompts constructed by first describing two co-occurring elements, and then replacing one with a less compatible alternative. This results in scenarios that are unlikely to occur in the real world.
Semantic Alignment in Text-to-Image Synthesis
@@
Text-to-image models often struggle to fully capture the semantic intent of input prompts, particularly when prompts involve complex or internally conflicting concepts. Previous works have analyzed common failure cases and proposed targeted improvements across various stages of the generation pipeline, including enhanced text embedding representations (Rassin et al., 2022; Feng et al., 2022; Tunanyan et al., 2023), refined attention mechanisms (Dahary et al., 2024) and guidance strategies that leverage attention maps for loss heuristics (Chefer et al., 2023; Rassin et al., 2023; Meral et al., 2024; Agarwal et al., 2023). Despite these advances, existing methods often fail to handle prompts containing contradictory concepts arising from the model’s learned associations. Our work directly addresses this underexplored challenge, focusing on contextually contradictory prompts.
Multi-Prompt Conditioning Techniques
@@
Conditioning diffusion models on multiple prompts has emerged as an effective strategy for improving control and compositionality. One line of work, primarily focused on personalization, introduces distinct learned tokens at different layers of the model and at various denoising timesteps (Alaluf et al., 2023; Voynov et al., 2023). This design allows each token to capture different attributes of the personalized concept, leading to improved identity preservation. Other approaches vary the prompt across timesteps to modulate specific visual properties, such as object shape (Liew et al., 2022; Patashnik et al., 2023), or alternate between rare and frequent object descriptions to improve attribute binding (Park et al., 2024). Fine-grained spatial control has also been achieved by assigning sub-prompts to separate image regions (Yang et al., 2024). Additionally, some methods leverage multiple diffusion models, each conditioned on different prompt attributes, and combine their outputs into a unified prediction (Liu et al., 2022; Bar-Tal et al., 2023). Unlike most prior works, we focus on utilizing multiple prompt to settle internal semantic tension, where concept combinations lead to contextual contradictions.
Figure 2. Our method navigates the denoising process using time-dependent proxy prompts. (Top) A large language model (LLM) is employed to decompose a target prompt with entangled or conflicting attributes into a sequence of proxy prompts and their corresponding time-step intervals. (Bottom) These intermediate prompts are injected into the diffusion process at the specified intervals. This time-dependent prompt conditioning aligns more closely with the model’s prior at each stage, enabling a semantically consistent and temporally coherent progression through the denoising process.
[…]
LLM-Guided Diffusion
@@
LLMs have demonstrated strong capabilities in language understanding. Moreover, they possess a broad, implicit grasp of world knowledge, acquired through large-scale training on diverse textual data. Recent approaches have leveraged these capabilities to guide diffusion model generation, often incorporating planning and reasoning to improve semantic alignment (Yang et al., 2024; Park et al., 2024; Hu et al., 2024). In our work, we employ LLMs with in-context learning to identify contextual contradictions, generate proxy prompts, and determine the corresponding timestep ranges for conditioning, while promoting a reasoning process through brief, explanatory outputs.
3. Method
@@
3.1. Coarse-to-Fine Denoising
As shown by previous works (Hertz et al., 2022; Balaji et al., 2022; Chefer et al., 2023; Patashnik et al., 2023), the denoising process of diffusion models operates in a coarse-to-fine manner: early steps determine low-frequency structures, while high-frequency details emerge in later steps.
This behavior is reflected in the training process, in which varying noise levels are applied across different timesteps, and the model is trained to reconstruct the original clean image from the noisy input. High noise levels used at early timesteps yield low signal-to-noise ratios, restricting reconstruction to spatially smooth, low-frequency components, such as background or layout, which are recoverable even when the image is heavily corrupted. At smaller timesteps, where the signal-to-noise ratio is higher, high-frequency features become more reliably reconstructible.
During inference, the noise is gradually removed as the model generates the image in a coarse-to-fine manner. This coarse-to-fine nature of the image generation leads to two key observations: (i) Irreversibility of details. At each stage of denoising, the model commits to a certain level of structural detail. Once a given attribute (e.g., layout, coarse shape) has been established, it becomes effectively fixed and cannot be revised in later steps, even if it misaligns with the prompt. (ii) Flexibility in high-frequency details. In early stages, high-frequency details have not yet emerged and thus are unaffected by the prompt. This allows greater flexibility in guiding the model without impacting the final generated fine details.
We demonstrate these observations in Figure 3, by examining the model’s $x_{0}$ predictions throughout the denoising process. In the top row, the prompt is fixed to “A howling wolf and a flying bat at midday”. In the middle row, we use the proxy prompt “A howling dog and a flying bird at midday” during the early steps of denoising, and then switch to the same target prompt used in the top row. We replaced “wolf” with “dog” because the two animals are semantically similar and share similar low-frequency visual features. Moreover, “a howling dog” is not contextually associated with nighttime. For similar reasons, we replaced “bat” with “bird”.
The first row illustrates the irreversibility of details. Night colors and a large moon are established early in the denoising process, and although they do not match the prompt, the model is unable to correct them in later steps.
The second column shows the flexibility of high-frequency details. Although different prompts were used for each row, the layout and object shapes are visually similar, reflecting the fact that distinguishing information between the target and proxy prompts – specifically regarding layout and object shape – has not yet emerged at this early stage of the denoising process. Yet, since a howling dog, unlike a howling wolf, is not entangled with night, the background in the second row adheres more closely to the prompt. By switching to the target prompt at this point in the process, we are still able to generate a wolf and a bat rather than a dog and a bird, while also aligning with the word “midday”.
The bottom row illustrates the importance of selecting an appropriate proxy prompt. Snakes and rats are associated with nighttime, which conflicts with the “midday” specification. In addition, their lack of visual similarity to the target objects leads to visual leakage between the bat and the wolf.
@@ 3.1. Coarse-to-Fine Denoising
These observations motivate a stage-aware prompting strategy that enables precise control over early structural decisions, while allowing more flexible conditioning for attributes that are not yet generated. Building on this, we now describe the two main components of our method: prompt decomposition and stage-aware prompt injection.
Figure 3. Coarse-to-fine denoising. $x_{0}$ predictions at the initial, intermediate, and final time steps. During the early denoising stages, where primarily low-frequency structures are generated, different prompts introducing distinct objects can produce visually similar layouts and objects. Once this structure is established, it becomes effectively frozen and cannot be modified in later steps. As shown in the bottom row, using incorrect placeholder objects in the proxy prompt can anchor erroneous visual information early on, leading to eventual misalignment to the target prompt.
3.2. Prompt Decomposition
@@
Given a prompt $P$ that contains contextually contradictory concepts, we aim to construct a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$ and corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$ that together reflect the intended semantics of $P$. Each proxy prompt $p_{i}$ is designed to preserve the relevant semantics of $P$ for attributes or concepts that typically form during its assigned interval $I_{i}$. This is crucial due to the irreversibility of details in the generation process. At the same time, $p_{i}$ is chosen to avoid contradictory combinations that tend to emerge at that stage, leveraging the flexibility in high-frequency details. This decomposition enables conditioning the diffusion model on contextually coherent content that evolves in tandem with the coarse-to-fine structure of the denoising process (see Figure 3).
To generate these proxy prompts and their associated intervals, we employ a large language model (LLM) with two complementary roles. First, we leverage the LLM’s broad world knowledge and reasoning capabilities to detect semantic contradictions and to propose contextually appropriate substitutes for conflicting concepts. Second, we use its instruction-following and in-context learning capabilities to infer the temporal staging of concepts, guided by examples and instructions that reflect the denoising behavior of diffusion models.
@@
We implement this using a structured prompt template that includes both instructional text and representative input-output examples, along with textual explanations of the contextual contradictions in the original prompts. These examples illustrate successful decompositions that resolve contextual conflicts, as well as cases that require no decomposition, enabling the LLM to generalize from both scenarios. The full instruction prompt is provided in the appendix (see Table 4).
In-context Examples
Our in-context examples for guiding the LLM include the target prompt as input and the corresponding proxy prompts, each paired with its associated timestep interval, as output. Each example also includes a brief explanation describing the contextual contradiction in the prompt, which is part of the expected output. By requiring the LLM to produce this explanation, we encourage it to engage in a reasoning process, improving its ability to identify contradictions and generate coherent proxy prompts. The examples cover a range of contradiction types, including semantic mismatches, spatial violations, behavioral implausibilities, and stylistic inconsistencies. We provide these examples in the appendix (see Table 5).
These examples were constructed by first identifying contextually-contradictory prompts that produce implausible generations in the base model (FLUX), particularly when concept conflicts appear early in the denoising process. Drawing on our coarse-to-fine analysis, we manually decomposed such prompts into proxy prompts sequences that stabilize low-frequency structure before gradually introducing high-frequency details. We used an LLM to generate natural language explanations and validate that each substitution preserve visual and contextual plausibility while resolving conflicts in a temporally coherent manner.
In total, we include 20 such examples, covering both effective decompositions and cases where minimal intervention was sufficient to preserve coherence. These examples guide the LLM in learning successful decomposition strategies, such as proxy substitution, temporal staging, and contradiction deferral.
Concept Substitution
A core mechanism of our prompt decomposition is concept substitution, which promotes contextual coherence by temporarily replacing a conflicting concept in the prompt with a structurally appropriate alternative (see Figure 3). A simpler alternative for avoiding contextual contradictions is to remove the conflicting concept entirely. However, we observe that introducing an object only at a later stage, without a placeholder proxy, can affect the perceived size of that object (see Figure 4). In extreme cases, introducing the object too late may result in it being omitted from the generated image altogether. Therefore, we prefer the concept substitution strategy.
As demonstrated in Figure 3, choosing the right substitution is key to the method’s success. When the substituted concept is semantically appropriate and visually similar to the original object during the interval in which it is applied, the model correctly associates it with the original concept. In contrast, poor substitutions – due to mismatches in scale, pose, location, or to inappropriate timing – can result in extra objects, layout errors, or failure to follow the prompt (see Figure 3, bottom row, and Figure 5).

### v2

# Image Generation from Contextually Contradictory Prompts

@@ Abstract
Figure 1: Our method addresses contextual contradictions in text-to-image generation. These contradictions arise when one concept implicitly conflicts with another due to the model’s learned associations. For example, if a concept like “butterfly” is strongly entangled with “flowers”, it may conflict with another concept in the prompt such as “bee’s hive”, leading the model to ignore or distort part of the semantic meaning.
1 Introduction
@@
Text-to-image diffusion models have demonstrated remarkable capabilities in generating high-quality and diverse visual content from natural language prompts [27, 24, 15]. However, achieving precise semantic alignment between the generated image and the conditioning prompt remains an open challenge. This issue becomes particularly pronounced when the target prompt lies outside the model’s training distribution, such as prompts that combine semantically plausible yet statistically uncommon or unconventional concepts. For example, as shown in Figure 1, generating an image from the prompt “A butterfly is in a bee’s hive” often results in a butterfly on a flower. This is due to the model’s prior that entangles butterflies with flowers, which implicitly contradicts the notion of a bee’s hive.
@@
We refer to this phenomenon as Contextual Contradiction, a conflict between two concepts that arises not from direct semantic opposition, but from the model’s associations learned during training. More precisely, we say that concept $A$ contextually contradicts concept $B$ if the model’s prior entangles $A$ with concept $C$, and $B$ contradicts $C$ (see Figure 2). In Figure 1, we illustrate this with a blowing dragon, which stands in contextual contradiction with the water due to its entanglement with fire.
Figure 2: On the left is a direct contradiction, since a girl cannot smile and frown at the same time. On the right is a contextual contradiction: while Disneyland and crying are not directly opposed, the model’s prior associates Disneyland with happiness, which conflicts with crying.
The phenomenon of contextual contradiction in text-to-image models relates to the broader issue of spurious correlations in deep learning. Models often exploit shortcuts, relying on correlations in the training data that are statistically strong but semantically misleading [13]. In this paper, we identify a similar bias in generative models: contextual contradictions occur when prompts combine concepts that individually align with the model’s priors but conflict when combined, revealing the model’s reliance on such correlations rather than robust compositional reasoning.
@@
To address this issue, we introduce Stage-Aware Prompting (SAP), which builds on the observation that the denoising process follows a coarse-to-fine progression, during which different semantic attributes (e.g., background, pose, shape, and texture) emerge at distinct stages [9, 4, 23]. Our key idea is to guide the model at each stage of denoising with the information most relevant to the type of content being formed at that point. To achieve this, we decompose the original prompt into a sequence of proxy prompts, each aligned with the attributes expected at a specific stage and designed to avoid contextual contradictions.
@@
Ensuring that proxy prompts preserve the original intent while avoiding contextual contradictions requires a broad understanding of the real world. For example, it involves understanding that a bear is entangled with specific poses, such as walking on all fours or standing upright, which in turn, contradicts the handstand pose. To achieve this, SAP leverages a large language model (LLM) to analyze the target prompt, identify contextual contradictions, and construct suitable proxy prompts.
@@
Through extensive experiments, we demonstrate the effectiveness of SAP in generating images from contextually contradictory text prompts. By introducing prompt information at targeted stages, SAP generates precise combinations of semantic attributes while avoiding undesired entanglement. Compared to previous methods, SAP’s stage-dependent prompt decomposition leads to more faithful and semantically aligned generations.
2 Related Work
Learned Spurious Correlations.
@@ 2 Related Work
Machine learning models are known to be sensitive to spurious correlations in their training data [13, 20, 33], leading to performance drops when training-time associations do not hold at test time. Prior work has extensively studied this in discriminative vision tasks, showing, for example, that recognition models tend to rely heavily on background cues [28, 31, 6]. In our work, we show that text-to-image models exhibit similar behavior. When given contextually contradictory prompts – combinations of concepts that conflict with correlations seen during training – diffusion models often fail to generate images that accurately reflect the prompt. We evaluate this behavior using the Whoops! dataset [7], which contains prompts constructed by first describing two co-occurring elements, and then replacing one with a less compatible alternative. This results in scenarios that are unlikely to occur in the real world.
Semantic Alignment in Text-to-Image Synthesis.
@@
Text-to-image models often struggle to fully capture the semantic intent of input prompts, particularly when prompts involve complex or internally conflicting concepts. Previous works have analyzed common failure cases and proposed targeted improvements across various stages of the generation pipeline, including enhanced text embedding representations [26, 12, 29], refined attention mechanisms [10], guidance strategies that leverage attention maps for loss heuristics [9, 25, 21, 2] and dynamic guidance scheduling via annealed classifier-free guidance [34]. Despite these advances, existing methods often fail to handle prompts containing contradictory concepts arising from the model’s learned associations. Our work directly addresses this underexplored challenge, focusing on contextually contradictory prompts.
“A howling wolf.” “A polar bear.” “A dragon blowing.”
Refer to caption Refer to caption Refer to caption
<howling> <polar> <dragon>
Refer to caption Refer to caption Refer to caption
(a) Entangled concepts.
“A woman writing.” “A woman writing with a dart.”
Refer to caption Refer to caption Refer to caption
<dart> <writing> log ratio
Refer to caption Refer to caption Refer to caption
(b) From entanglement to contradiction.
Figure 3: (a) By examining attention maps, we observe that textual tokens embed contextual associations, leading to the generation of concepts not explicitly mentioned in the prompt. For example, the token ‘howling’ encourages the presence of the moon, as indicated by its strong attention connection. (b) In prompts with contextual contradictions, two tokens may overlap in attending to the same region, as seen with ‘dart’ and ‘writing’. The token ‘writing’ dominates this region, as shown in the log-ratio map, where red areas indicate stronger attention to ‘writing’ relative to ‘dart’.
Multi-Prompt Conditioning Techniques.
@@
Conditioning diffusion models on multiple prompts has emerged as an effective strategy for improving control and compositionality. One line of work, primarily focused on personalization, introduces distinct learned tokens at different layers of the model and at various denoising timesteps [3, 30]. This design allows each token to capture different attributes of the personalized concept, leading to improved identity preservation. Other approaches vary the prompt across timesteps to modulate specific visual properties, such as object shape [18, 23], or alternate between rare and frequent object descriptions to improve attribute binding [22]. Fine-grained spatial control has also been achieved by assigning sub-prompts to separate image regions [32]. Additionally, some methods leverage multiple diffusion models, each conditioned on different prompt attributes, and combine their outputs into a unified prediction [19, 5]. Unlike most prior works, we focus on utilizing multiple prompts to settle internal semantic tension, where concept combinations lead to contextual contradictions.
LLM-Guided Diffusion.
@@
LLMs have demonstrated strong capabilities in language understanding. They also capture broad world knowledge through large-scale training on diverse text. Recent approaches have leveraged these capabilities to guide diffusion model generation, often incorporating planning and reasoning to improve semantic alignment [32, 22, 16]. In our work, we employ LLMs with in-context learning to identify contextual contradictions, generate proxy prompts, and determine the corresponding timestep ranges for conditioning, while encouraging reasoning through brief explanatory outputs.
3 From Entangled Concepts to Contextual Contradictions
Diffusion models inherit strong distributional biases from their training data, where objects are frequently tied to specific contexts. For example, prompts like “a duck” almost always result in water backgrounds, and “a polar bear” appears in snow. These reflect entangled concepts, learned associations that go beyond the explicit text. While often helpful, such priors hinder generation when prompts require unusual or contradictory combinations.
To study this effect, we analyze attention maps (Figure 3(a)), which indicate the spatial regions influenced (attended) by each token. We find that text tokens often attend not only to the image regions directly corresponding to the object but also to contextually linked elements. For example, in “a howling wolf”, the token ‘howling’ influences both the mouth and the moon. Similarly, ‘dragon’ attends to flames even when fire is not mentioned. These patterns reveal that the model encodes distributional correlations beyond literal semantics.
These distributional correlations contribute to the difficulty of generating images with contextual contradictions. For example, when generating an image from the prompt “a woman writing with a dart”, the model fails to replace the pen with a dart. The attention maps shed light on this failure: both ‘writing’ and ‘dart’ attend to the same spatial area (the hand/tool), but ‘writing’ dominates (see log-ratio maps in Figure 3(b)), suppressing the influence of ‘dart’. This reflects a failure mode in which entrenched associations override less familiar ones, preventing proper integration of conflicting concepts.
4 Stage Aware Prompting (SAP)
Figure 4: Coarse-to-fine denoising with stage-aware prompting. We show $x_{0}$ predictions at initial, intermediate, and final steps. A is the target prompt, B a suitable proxy, and C an unsuitable proxy. Using A alone locks in night and moon despite “midday”. Using B first and then switching to A preserves a daytime layout and later adapts the identities to wolf and bat. Using C first sets a layout without a flying object, so the final image fails to produce the intended subjects.
@@ 4 Stage Aware Prompting (SAP)
4.1 Coarse-to-Fine Denoising
Diffusion models generate images in a coarse-to-fine manner: early steps determine low-frequency structures, while high-frequency details emerge in later steps [14, 4, 9, 23]. From this behavior, we draw two key observations: (i) Irreversibility of details. As denoising progresses, the model sequentially commits to different levels of detail, beginning with layout and overall shape. Once these are formed, they cannot be revised in later stages, even if they conflict with the prompt. (ii) Flexibility in high-frequency details. In early stages, high-frequency details are absent and unaffected by the prompt, enabling flexible guidance without influence on fine details.
As shown in Section 3, contextual contradictions stem from concept entanglement in the diffusion prior. Since different concepts emerge at different levels of detail during the coarse-to-fine denoising process, they can be decomposed across the denoising stages. We illustrate this in Figure 4 by examining the model’s $x_{0}$ predictions across denoising steps. In the top-left row, the prompt “a howling wolf and a flying bat at midday” shows that early steps already impose entangled nighttime and moon structures, contradicting “midday”. In the bottom-left row, starting with a proxy prompt containing “dog” and “bird” (instead of “wolf” and “bat”, respectively) and later switching to the target prompt produces a correct midday scene with the intended objects. The bottom-left row demonstrates both the irreversibility of coarse details (the scene remains daytime) and the flexibility of high-frequency details (the object identities adapt). The top-right row highlights the importance of selecting an appropriate proxy prompt. Since rats do not fly, the layout determined in early steps lacks a flying object, resulting in a sitting wolf with bat wings.
@@ 4.1 Coarse-to-Fine Denoising
These observations motivate a stage-aware prompting strategy that fixes structural decisions early while keeping later attributes flexible. Building on this, we now describe the two main components of our method: prompt decomposition and stage-aware prompt injection.
4.2 Prompt Decomposition
Figure 5: SAP generates images from contextually contradictory prompts using time-dependent proxy prompts. (Top) A large language model (LLM) decomposes the target prompt into a sequence of proxy prompts with corresponding timestep intervals. (Bottom) These proxy prompts are injected into the diffusion process at their designated intervals to guide generation.
@@ 4.2 Prompt Decomposition
Given a prompt $P$ that contains contextually contradictory concepts, we aim to construct a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$ and corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$ that together reflect the intended semantics of $P$. Each proxy prompt $p_{i}$ is designed to (i) preserve the relevant semantics of $P$ for attributes typically formed during its interval $I_{i}$, which is crucial due to the irreversibility of details, and (ii) avoid contradictions likely to emerge at that stage, leveraging the flexibility in high-frequency details. This decomposition conditions the diffusion model on contextually coherent content that evolves in tandem with the coarse-to-fine denoising process.
@@
To generate proxy prompts and their intervals, we use a large language model (LLM) that detects contextual contradictions and proposes suitable substitutes for conflicting concepts. It also infers the appropriate staging of these concepts within the proxy prompts. We implement this using a structured prompt template containing instructional text, in-context examples, and explanations of contextual contradictions. The examples demonstrate both successful decompositions and cases requiring no decomposition, enabling the LLM to generalize. The full instruction prompt is provided in the Appendix.
In-context Examples.
Our in-context examples take a target prompt as input and output proxy prompts with timestep intervals, along with a brief explanation of the contradiction. Requiring the LLM to provide this explanation encourages reasoning about conflicts and ensures coherent substitutions. These examples were created by identifying contextually contradictory prompts that fail under the base model (FLUX) and manually decomposing them into proxy prompts with corresponding intervals. The explanations were auto-generated by an LLM. We include 20 examples that demonstrate diverse strategies for handling contradictions. We next elaborate on one of the most frequent strategies in our method.
Concept Substitution.
In this strategy, a conflicting concept is temporarily replaced with a structurally appropriate proxy (Figure 4). A simpler alternative is to omit the conflicting concept, but introducing an object only in later stages without a placeholder can distort its perceived size or cause it to be omitted entirely. In Figure 6, we demonstrate this by comparing decompositions that differ only in when the second interval begins. The first proxy specifies the background, while the second adds the foreground. Introducing the foreground early allows the model to allocate more space, whereas delaying it constrains the layout and produces a smaller object. Substitution resolves these issues and yields stable layouts. In Figure 7, we show the effect of misplacing intervals across denoising stages. Using two proxies with different intervals, where the second is the full prompt, we observe two failure modes: introducing the full prompt too early prevents disentangling contradictions, while introducing it too late alters only fine details. Earlier in Figure 4 we illustrated the importance of careful proxy selection.
4.3 Stage-aware Prompt Injection
Given a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$, their corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$, and a text-to-image (T2I) diffusion model, we condition the model using different prompts throughout the denoising process. At each timestep $t$, we apply the prompt $p_{i}$ such that $t\in I_{i}$. By aligning each proxy prompt with its interval, the denoising process is guided by concepts appropriate to the level of detail emerging at that stage. This enables gradual image construction while avoiding conflicts with the model’s learned priors. The injection mechanism integrates seamlessly into existing inference pipelines without architectural modifications and is compatible with a range of pretrained diffusion models.
5 Experiments and Results
In this section, we evaluate SAP through qualitative (Section 5.1) and quantitative (Section 5.2) comparisons. We further conduct ablation studies (Section 5.3) on component contributions and robustness, followed by a discussion on the limitations of our method.